OpenStudy (anonymous):

what is the perimeter of a right triangle with the legs of 15cm and 20cm

OpenStudy (anonymous):

use pythagorean theorem to solve for the hypotenuse. c^2 = 15^2 + 20^2. c is going to be 25.

OpenStudy (anonymous):

thus the answer is 15+20+25 = 60cm.

OpenStudy (anonymous):

so the perimeter is 60cm?

OpenStudy (anonymous):

easier way to solve it is that you recognize that it's a pythagorean triple. it's similar to a 3-4-5 triangle, so you know that its going to be 15-20-25 (multiplied by 5)

OpenStudy (anonymous):

yes. perimeter is 60cm

OpenStudy (anonymous):

okay thanks